Hult Games buys electronic components for manufacturing from two suppliers, Milan Components and Dundee Parts. If the components are delivered late, the shipment to the customer is delayed. Delayed shipments lead to contractual penalties that call for Hult to reimburse a portion of the purchase price to the customer.

During the past quarter, the purchasing and delivery data for the two suppliers showed the following.

Milan Dundee Total
Total purchases (cartons) 147,000 63,000 210,000
Average purchase price (per carton) $ 10.00 $ 12 $ 10.60
Number of deliveries 80 20 100
Percentage of late deliveries 25 % 10 % 22 %

The Accounting Department recorded $861,000 as the cost of late deliveries to customers.

Assume that the average quality, measured by the percentage of late deliveries, and prices from the two companies will continue as in the past.

Required:

What percentage of late deliveries by Milan would make Hult indifferent between buying from Milan Components and Dundee Parts?
Answer #1

1) we will find cost of late delivery per cartoon.

number of cartoon delivered late = (147000*25%)+(63000*10%)

=36750+6300

=43050 cartoons.

cost per cartoon = 861000$/43050

=20$ per cartoon.

2) cost per cartoon of Dundee

Dundee
purchase cost 12
percentage of late deliveries 10%
cost per cartoon for late delivered cartoon 20$
equivalent cost of late delivery per cartoon [20$*10%] $2
Equivalent cost oer cartoon [12+2] $14

3) percentage of late deliveries by Milan would make Hult indifferent between buying from Milan Components and Dundee Parts. Suppose X%

cost of milan = cost of Dundee

$10 + [$20cost per late delivery *X%] = $14

$20*X% =$14-$10

X% = $4/$20

X =20%

Answer 20%.